OpenStudy (anonymous):

The legs of a right triangle are 9 cm and 40 cm long. What is the length of the hypotenuse? 41cm 45cm 49cm 50cm

OpenStudy (lalaly):

41

OpenStudy (anonymous):

41

OpenStudy (anonymous):

41

OpenStudy (anonymous):

hypotenuse^2=9^2+40^2 =81+1600 =1681 hypotenuse=sqrt1681 =41